### v1

Abstract
The standard mixture modelling framework has been widely used to study heterogeneous populations, by modelling them as being composed of a finite number of homogeneous sub-populations. However, the standard mixture model assumes that each data point belongs to one and only one mixture component, or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic. It is in fact conceptually very different to represent an observation as partly belonging to multiple groups instead of belonging to one group with uncertainty. For this purpose, various soft clustering approaches, or individual-level mixture models, have been developed. In this context, Heller et al (2008) formulated the Bayesian partial membership model (PM) as an alternative structure for individual-level mixtures, which also captures partial membership in the form of attribute-specific mixtures, but does not assume a factorization over attributes.
Our work proposes using the PM for soft clustering of count data arising in football performance analysis and compares the results with those achieved with the mixed membership model and finite mixture model. Learning and inference are carried out using Markov chain Monte Carlo methods.
The method is applied on Serie A football player data from the 2022/2023 football season, to estimate the positions on the field where the players tend to play, in addition to their primary position, based on their playing style. The application of partial membership model to football data could have practical implications for coaches, talent scouts, team managers and analysts. These stakeholders can utilize the findings to make informed decisions related to team strategy, talent acquisition, and statistical research, ultimately enhancing performance and understanding in the field of football.
keywords:
Partial membership models, Model based clustering, Finite mixture models, Sports data analysis, Football analytics.
1 Introduction
Model-based clustering has been widely used among researchers to study heterogeneous populations, by modelling them as being composed of a finite number of homogeneous sub-populations (eg. Fraley and Raftery, 2002; McLachlan and Peel, 2000; Bouveyron et al, 2019; Gormley et al, 2023). In this framework, observations within a dataset are modeled as originating from one of multiple probability distributions. The objective is to achieve a clustering solution where observations are partitioned into distinct groups. Notably, observations that show a significant posterior probability of belonging to more than one cluster are considered to have uncertain group membership. This uncertainty can sometimes suggest a model that does not adequately fit the data. However, the standard mixture model assumes that each data point belongs to one and only one mixture component, or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic.
In contrast, mixed membership and partial membership models accommodate partial membership to multiple clusters. For example, let’s consider a football player with the role of a trequartista. His duties on the pitch change depending on the phase of play. In an offensive phase, they are those associated with the attacking role, while in a defensive phase they are those associated with the midfielder role. That football player should be represented as partly belonging to two different classes or sets. Being certain that a player’s role is partly midfielder and partly striker is very different from being uncertain about a player’s role on the pitch.
The foundational concept of mixed membership modeling traces its origins to the 1970s with the development of the Grade of Membership (GoM) model by mathematician Max Woodbury, which was designed for “fuzzy” classifications in medical diagnostic scenarios (Woodbury et al, 1978). It was not until the early 2000s, following the widespread adoption of Bayesian methods and an enhanced understanding of the duality between the discrete and continuous nature of latent structures in the GoM model, that a new Bayesian approach to the GoM model was developed. Independently, and within a short span of time, three mixed membership models emerged to address challenges in distinctly different domains: Blei et al (2003) – Latent Dirichlet Allocation (LDA); Erosheva (2003) – Grade of Membership model (GoM); Pritchard et al (2000) – Admixture model. Mixed membership models unify the LDA, GoM, and admixture models in a common framework and provides ways to construct other individual-level mixture models by varying assumptions on the population, sampling unit and latent variable levels, and the sampling scheme.
Partial membership models (Heller et al, 2008) are defined using a similar framework, but they overcome some of the drawbacks of mixed membership models. In the present paper we specify partial membership models for count data, in particular when component distributions are Poisson distributions.
Eventhough partial membership models exist for some time, they are still under utilised in literature. Gruhl and Erosheva (2013) apply partial membership model to NBA player data from the 2010–11 season and compare the results with those achieved with the mixed membership model. Chen et al (2017) used partial membership for semantic image segmentation, while Hou-Liu and Browne (2022) start from partial membership models and epistatic clustering (Zhang, 2013) to develop an “hybrid method” between the two, to cluster hybrid species of flowers, which tend to exhibit a mixture of parent characteristics.
We apply the PM model to Serie A football player performance data, from the 2022/2023 football season. The aim is to estimate the various roles players are inclined to occupy on the field, beyond their primary positions, by analyzing their playing styles. The goal of the application within sports analysis is to assist coaches, talent scouts, team managers, and analysts in making more informed decisions. These decisions pertain to team strategy, talent acquisition, and advancing statistical research in sports.
Mixed membership model for count data, outlined in White and Murphy (2016), and mixture model, are also applied and the results are compared. The comparison suggests that in this application, the partial membership model gives more realistic and interpretable results.
The article is organised as follows. Section 2 outlines the Bayesian partial membership model specification with Poisson component distributions, and compares the data generative process to those in mixed membership and mixture models. It also gives an overview on technical aspects like label switching and model selection, justifying the choice of the information criterion through literature and a simulation. In Section 3, both partial and mixed membership models are applied to Serie A football players performance data and the results are compared. For sake of completeness, a Poisson mixture model is also applied to this study. However, the results are only briefly explored, because a crisp clustering approach does not align with the specific objectives and purposes of the proposed application. Conclusions and future developments follow in Section 4. To further demonstrate the versatility and effectiveness of the proposed model, we include an additional application of the model to Washington DC bike sharing data in Appendix A.
2 Partial Membership Model
2.1 Mixture Model
Consider a data set $\bm{X}=\{\bm{x}_{ij}:i=1,2,\dots,N,j=1,2,\dots,J\}$ with $N$ observations and $J$ features per observation. In a finite mixture model the data are modeled as a mixture of $K$ component densities $P_{k}(\cdot\mid\bm{\theta}_{k})$ with unknown mixing proportions ${\pi}_{1},\dots,{\pi}_{K}$. The density of $\bm{x}_{i}$ is given as:
$P(\bm{x}_{i}\mid\bm{\Theta},\bm{\pi})=\sum_{k=1}^{K}{\pi}_{k}P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k}),$
where $\bm{\pi}=\{{\pi}_{k}:k=1,2,\dots,K\}$ are the mixing proportions and $\bm{\Theta}=\{\bm{\theta}_{k}:k=1,2,\ldots,K\}$ are the mixture component parameters. In this framework, a data point is assumed to come from one (and only one) of the $K$ mixture components. Let $\bm{\tau}_{i}$ be the latent component membership indicator variable for observation $i$, so $\tau_{ik}\in\{0,1\}$ and $\sum_{k}\tau_{ik}=1$. In the mixture model, we assume that $\bm{\tau}_{i}\sim\mbox{multinomial}(1,\bm{\pi})$, and so $\bm{\tau}_{i}=(\tau_{i1},\dots,\tau_{iK})$ is the binary membership vector, where $\tau_{ik}=1$ indicates that the data point $\bm{x}_{i}$ was generated by mixture component $k$.
2.2 Bayesian Partial Membership Model
In a partial membership model, we relax the constraint $\tau_{ik}\in\{0,1\}$ to take any continuous value in the range $[0,1]$ and where $\sum_{k}\tau_{ik}=1$; therefore, $\tau_{ik}$ is in the unit simplex, allowing multiple cluster memberships for each data point. In the partial membership model, the density of an observation becomes,
$P(\bm{x}_{i}\mid\bm{\Theta})\propto\int_{\bm{\tau}_{i}}P(\bm{\tau}_{i})\prod_{k=1}^{K}P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k})^{\tau_{ik}}d\bm{\tau}_{i}.$
We integrate over all values of $\bm{\tau}_{i}$ instead of summing. The product over $k$ reflects that the contribution to the likelihood of $\bm{x}_{i}$ from each component is compounded multiplicatively, representing how each component contributes to explaining $\bm{x}_{i}$ given its partial membership weights $\tau_{ik}$. The exponent $\tau_{ik}$ on $P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k})$ represents the degree of membership of $\bm{x}_{i}$ in component $k$. This is not directly present in the finite mixture model, where the component membership is implicit in the mixing weights $\pi_{k}$ and assumed to be binary (fully in one component). In the partial membership context, $\tau_{ik}$ as an exponent softens this assumption, allowing $\bm{x}_{i}$ to “partially belong” to multiple components simultaneously, each to varying degrees expressed by $\tau_{ik}$.
In this work, we focus on the case when the form of the distribution for each cluster $P_{k}(\bm{x}_{i}\mid\bm{\theta}_{i})$ are independent Poisson distributions:
$P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k})=\prod_{j=1}^{J}P_{k}(\bm{x}_{ij}\mid\lambda_{kj})=\prod_{j=1}^{J}\frac{\lambda_{kj}^{\bm{x}_{ij}}e^{-\lambda_{kj}}}{\bm{x}_{ij}!},$
where $\bm{\theta}_{k}=(\lambda_{k1},\lambda_{k2},\ldots,\lambda_{kJ})$. Thus, the partial membership model assumes that each data point is drawn from
$x_{ij}\mid\bm{\tau}_{i}\sim\mbox{Poisson}\left[\exp\left(\sum_{k=1}^{K}\tau_{ik}\log\lambda_{kj}\right)\right].$
Furthermore, we assume that the partial membership weights are drawn from a Dirichlet distribution:
$\bm{\tau}_{i}\sim\mbox{Dirichlet}(\bm{\delta}).$
We consider a model with $K$ clusters and specify prior distributions for the model parameters. In particular, we let $\bm{\delta}$ be a $K$-dimensional vector of positive hyperparameters $\bm{\delta}\sim\mbox{unif}(a,b)$; the choice of the Uniform distribution for $\delta$ and the values of $(a,b)$ will be explained in Section 2.5. We use a gamma conjugate prior distribution for the Poisson parameters
$P(\lambda_{kj})\propto\lambda_{kj}^{\alpha-1}e^{-\beta\lambda_{kj}},$
where $\alpha$ and $\beta$ are hyperparameters of the prior. We found that the use of priors with different (sensible) choices of hyperparameters were found to have little effect on the clustering obtained in Section 3 and Appendix A. A graphical model representation of the partial membership model is given in Figure 1.
Figure 1: Graphical model of the partial membership formulation
2.3 Mixed Membership Model
A method for fitting mixed membership models to count data is outlined in White and Murphy (2016). Broadly speaking, mixed membership (MM) models operate under the assumption that each data feature (such as pixels in an image analysis scenario) of a given item (e.g., an image) is independently drawn from a mixture distribution defined by the membership vector of the item, $x_{ij}\mid\bm{\tau}_{i}\sim\sum_{k}\tau_{ik}P({x}_{ij}\mid\lambda_{kj})$ for $j=1,2,\ldots,J$. MM models make sense when the items being analyzed, such as images, can be seen as collections of exchangeable parts (e.g., pixels). In contrast, partial membership models do not rely on the assumption of exchangeable sub-objects.
The generative process for $\bm{X}$ in the partial membership model, compared to those in mixed membership and mixture models in Table 1.
Table 1: Comparison of the data generation process for the mixture model, mixed membership model and partial membership models for count data.
Mixture Mixed Membership Partial Membership
for($i$ in $1:N$) for($i$ in $1:N$) for($i$ in $1:N$)
  $\bm{\tau}_{i}\sim\mbox{multinomial}(1,\bm{\pi})$   $\bm{\tau}_{i}\sim\mbox{Dirichlet}(\bm{\delta})$   $\bm{\tau}_{i}\sim\mbox{Dirichlet}(\bm{\delta})$
  for($j$ in $1:J$)   for($j$ in $1:J$)   for($j$ in $1:J$)
    $\bm{Z}_{ij}\sim\mbox{multinomial}(1,\bm{\tau}_{i})$     $\mu_{ij}=\exp(\sum_{k=1}^{K}\tau_{ik}\log\lambda_{kj})$
    $\bm{X}_{ij}\sim\mbox{Poisson}(\lambda_{\bm{\tau}_{i},j})$     $\bm{X}_{ij}\sim\mbox{Poisson}(\lambda_{Z_{ij},j})$     $\bm{X}_{ij}\sim\mbox{Poisson}(\mu_{ij})$
2.4 Inference
The PM posterior distribution takes the form:
$P(\bm{\tau},\bm{\lambda},\bm{\delta}\mid\bm{x},\alpha,\beta,a,b)$ $\propto$ $\prod_{i=1}^{n}\left[P(\bm{\tau}_{i}\mid\bm{\delta})\prod_{k=1}^{K}\prod_{j=1}^{J}P_{k}(\bm{x}_{ij}\mid\lambda_{kj})^{\tau_{ik}}\right]$
×$\times P(\bm{\lambda}\mid\alpha,\beta)P(\bm{\delta}\mid a,b)$
Fitting the PM consists of inferring all unknown variables given $\bm{X}$, for which we employ a Bayesian approach using Monte Carlo Markov chain (MCMC). A notable advantage of PM over Mixed Membership (MM) models is that MM models require a discrete latent variable for each sub-object to indicate the mixture component from which it was drawn. This substantial number of discrete latent variables can make MCMC sampling in MM models significantly more challenging than in PM models.
The model was fitted using NIMBLE, a versatile framework for developing statistical algorithms for general model structures within R (de Valpine et al, 2017; de Valpine et al, 2023). NIMBLE facilitates the definition of models as directed acyclic graphs (DAGs), which simplifies the deployment of a range of algorithms. It accommodates several MCMC techniques and provides a comprehensive array of resources for model evaluation, selection, and prediction error assessment. These features led us to prefer NIMBLE over other Bayesian MCMC software like Stan and JAGS.
2.5 Identifiability
Parameters in a model are not identified if the same likelihood function is obtained for more than one choice of the model parameters (Teicher, 1963; Gelman et al, 2013).
The issue of identifiability in finite mixture distributions demands careful consideration, particularly when conducting Bayesian inference. Frühwirth-Schnatter (2006) describes three types of non-identifiability: invariance to relabeling of the components in the mixture model, non-identifiability due to potential overfitting, and a generic non-identifiability in certain classes of mixture distributions. Non identifiability can lead to poor convergence of MCMC methods used in Bayesian inference.
To draw valid posterior inference, the invariance to relabeling characteristic requires particular attention. Among the possible solutions, such as Celeux et al (2000) or Murphy et al (2020), in this paper the label switching problem has been addressed permuting after the run of the model the labels to each MCMC draw, using the probabilistic relabelling algorithm of Sperrin et al (2010), provided by the R package $\mathsf{label.switching}$.
In the application proposed in the present paper, potential identifiability challenges are addressed through the incorporation of entities termed as archetypal units. These units emerge from a statistical technique known as archetypal analysis. The primary objective of archetypal analysis is to identify a limited number of extreme, yet possibly hypothetical, data points (referred to as the archetypal units) within a multivariate dataset. The essence of this method is that all observed data points can be effectively represented as convex combinations of these archetypes, which form what it is called the “performance profile”. These archetypes are themselves constrained to be convex combinations of actual data points and lie on the data set boundary, i.e., the convex hull. This statistical method was first introduced by Cutler and Breiman (1994) and has found applications in different areas, e.g., in economics (Porzio et al, 2008), astrophysics (Chan et al, 2003), pattern recognition (Bauckhage and Thurau, 2009) and sports analytics (Eugster, 2012; Seth and Eugster, 2016).
In addressing the identifiability issue within the PM model, we employed archetypal analysis, guided by the presence of units who distinctly belong to a single cluster. These units, who exhibit clear archetype-defining characteristics of a particular cluster, provide anchor points. These anchors not only crystallize the definition of each cluster but also enhance the overall stability and interpretability of the model. They serve as concrete examples against which the characteristics of more ambiguously classified units can be compared, offering a benchmark for cluster assignment. Moreover, the presence of such well-defined units aids in delineating the boundaries of each cluster. This clarity is crucial in a partial membership context, where the risk of overly fluid cluster definitions could lead to diminished utility of the model. By having these distinctly categorized units, namely football players in the proposed application, we maintain a balance in our model; it remains flexible enough to accommodate the partial memberships of players who exhibit mixed attributes, yet it is anchored firmly enough to prevent the clusters from becoming indistinct or overlapping excessively.
We choose a uniform distribution with parameters $(0,10)$ as the hyperprior for the $\bm{\delta}$ parameter in the mixture weights from a Dirichlet distribution: $\tau_{i}\sim\mbox{Dirichlet}(\bm{\delta})$. This decision was instrumental in ensuring the identification of archetypal players, while preserving the interpretability of the results presented in Section 3. It is important to note that for sake of completeness, we experimented with increasing the sparsity of the mixture weights matrix by opting for a smaller scalar parameter for the Dirichlet distribution (Wang and Blei, 2009). However, this increased sparsity led to a higher optimal number of clusters according to the information criterion used. Consequently, the uniform distribution emerged as our preferred choice. It effectively facilitated the identification of archetypal units while maintaining a manageable number of clusters, thus ensuring the results remained interpretable.
2.6 Model Selection
According to Watanabe and Opper (2010), a statistical model is said to be regular if the map taking parameters to probability distributions is one-to-one and if its Fisher information matrix is positive definite. If a model is not regular, then it is said to be singular. If a statistical model contains a hierarchical structure or latent variables then the model is generally singular. In singular statistical models, the maximum likelihood estimator does not satisfy asymptotic normality. Consequently, standard model selection criteria like the Akaike Information Criterion (AIC) and the Bayes Information Criterion (BIC) are not always appropriate for model selection. WAIC (Widely Applicable Information Criterion) (Watanabe and Opper, 2010) can be used for estimating the predictive loss of Bayesian models, using a sample from the full-data posterior, and it is applicable to non-regular models such as the PM.
In the present paper, WAIC is calculated from Equations 5, 12, and 13 in Gelman et al (2014), and it is the log pointwise predictive density minus a correction for effective number of parameters to adjust for overfitting. According to Millar (2018), the marginalized WAIC might be more accurate for choosing the correct model.
We fit the model in a simulated data scenario to verify the number of times the correct model is selected by the WAIC conditional on all of the parameters (WAICc) and marginalized over $\bm{\tau}$ (WAICm). We generated 100 random membership vectors from a Dirichlet($\bm{\delta}$) distribution with shape parameter $\bm{\delta}$ = (0.5, 0.5, 0.5, 0.5). Using these membership scores, we simulated 100 partial membership models with $N=200$, $J=25$ and $K=4$ to match the football players application scenario. We ran the MCMC algorithm for 20000 iterations, keeping every 50th draw. We discarded the first 5000 iterations as burn-in. MCMC was run the model over a range of values of $K=1,...,8$. To assess convergence, we examined trace plots.
The simulation results are illustrated in Table 2 and the simulation confirms that the marginalized WAIC, is more accurate for choosing the right model, with $99\%$ of success.
Table 2: Number of times each $K$ corresponds to the minimum WAIC conditional to all the parameters (WAICc) and marginalized for $\bm{\tau}$ (WAICm). $N=200$, $J=25$, Number of runs=100, true $K=4$
WAICc WAICm
$K=1$ 0 -
$K=2$ 0 0
$K=3$ 0 0
$K=4$ 79 99
$K=5$ 16 1
$K=6$ 3 0
$K=7$ 2 0
$K=8$ 0 0
3 Serie A Football Players
We analyze the performance statistics of the 192 Serie A football players who played more than 1350 minutes during the 2022/2023 football season [^1] The data are freely available at https://fbref.com.. The analysis considers a set of 22 count variables for each player across the games that they played, selected to encompass the essential skills associated with each player’s role on the field.
This application allows us to verify the reliability of the model’s results by comparing them with each player’s designated playing position. The player’s partial membership allows us to estimate the positions on the field where the players tend to play, in addition to their (primary) designated position, based on their playing style.
In the field of clustering football players’ positions, a study with objectives similar to ours was conducted by Seth and Eugster (2016), employing archetypal analysis. A key distinction between our study and that study lies in the methodology adopted, as well as the nature of the data used. Seth and Eugster (2016) based their analysis on skill ratings from the PES Stats Database (https://www.pesmaster.com), a community-driven platform that compiles statistics and skill ratings for soccer players. This database was originally developed for the video game “Pro Evolution Soccer” by Konami. In contrast, our study utilizes actual statistical data recorded during games, providing a different perspective and basis for analysis.
The performance variables included in the analysis and modelling are given as follows.
  • Gls – Number of goals.
  • Ast – Number of assists.
  • PrgC – Progressive carries: carries that move the ball towards the opponent’s goal line at least 10 yards from its furthest point in the last six passes, or any carry into the penalty area. Excludes carries which end in the defending 50% of the pitch.
  • PrgP – Progressive Passes: progressive Passes completed passes that move the ball towards the opponent’s goal line at least 10 yards from its furthest point in the last six passes, or any completed pass into the penalty area. Excludes passes from the defending 40% of the pitch.
  • Sh – Shots Total: does not include penalty kicks.
  • SoT – Shots on Target: does not include penalty kicks.
  • KP – Key Passes: Passes that directly lead to a shot (assisted shots).
  • PiFT – Passes into Final Third: completed passes that enter the 1/3 of the pitch closest to the goal, not including set pieces.
  • PPA – Passes into Penalty Area: completed passes into the 18-yard box, not including set pieces.
  • CrsPA – Crosses into Penalty Area: completed crosses into the 18-yard box, not including set pieces.
  • SCA – Shot-Creating Actions: the two offensive actions directly leading to a shot, such as passes, take-ons and drawing fouls.
  • PassLive – SCA (PassLive): completed live-ball passes that lead to a shot attempt.
  • PassDead – SCA (PassDead): completed dead-ball passes that lead to a shot attempt. Includes free kicks, corner kicks, kick offs, throw-ins and goal kicks.
  • TO – SCA (TO): successful take-ons that lead to a shot attempt.
  • ShToSh – SCA (Sh): shots that lead to another shot attempt.
  • Def – SCA (Def): defensive actions that lead to a shot attempt.
  • GCA – Goal-Creating Actions: the two offensive actions directly leading to a goal, such as passes, take-ons and drawing fouls. Note: a single player can receive credit for multiple actions and the shot-taker can also receive credit.
  • Tkl – Tackles: Number of players tackled
  • Blocks – Number of times blocking the ball by standing in its path
  • Int – Interceptions
  • Clr – Clearances
  • Err – Errors: Mistakes leading to an opponent’s shot
Partial membership (PM), mixed membership (MM) and mixture models are applied to the dataset, with the WAIC suggesting that the $K=4$, $K=5$ and $K=6$ profile models are optimal, as illustrated in Table 3.
Table 3: WAIC values for the partial membership, mixed membership and mixture models, with $K$ ranging between 2 and 8. The optimal model for each model type is highlighted in boldface.
WAIC $K=2$ $K=3$ $K=4$ $K=5$
PM 52237.45 44274.20 37692.89 47919.76
MM 35330.92 29833.80 28309.35 27898.17
Mixture 35385.72 29846.14 28369.55 27981.13
$K=6$ $K=7$ $K=8$
PM 42843.47 43603.61 41295.05
MM 28096.84 28480.93 29099.70
Mixture 27900.42 28219.43 28242.52
3.1 Partial Membership Model Results
Table 4 and Figure 2 give the mean profiles for the partial membership model with $K=4$.
Figure 2: PM model. Expected profiles means, conditional on profile membership, with 4 profiles.
Table 4: PM model expected profiles means
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 0.00 0.04 0.13 0.32 0.00 0.00 0.13 8.02
2 0.65 0.19 6.42 52.34 8.39 2.11 2.21 45.67
3 2.74 3.75 71.57 170.20 29.29 9.25 45.91 136.88
4 15.14 3.70 42.25 19.69 88.57 34.20 26.75 6.75
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 0.19 0.11 1.45 1.09 0.77 0.01 0.00 0.07
2 1.79 0.36 8.18 7.68 0.04 0.11 0.95 0.63
3 41.41 15.58 87.28 63.60 22.84 2.82 3.11 1.05
4 14.34 2.34 56.61 34.81 1.66 10.87 10.39 0.48
GCA Tkl Blocks Int Clr Err
1 0.09 0.36 0.06 0.26 6.87 1.34
2 0.59 39.94 34.80 34.11 119.34 1.01
3 8.64 56.30 29.47 24.78 25.48 0.43
4 7.48 7.45 8.20 2.13 6.13 0.10
The profile means can be interpreted in the following way:
  • Profile 1 consistently exhibits low values compared to the other profiles across almost every variable, except for Err – Errors. This profile can be interpreted as grouping the goalkeepers, as they typically have significantly lower average values in the considered variables. These variables are more relevant to players in more active roles involving ball possession.
  • Profile 2 is characterized by notably high values in Clr – Clearances, Int – Interceptions, Blocks – Number of times blocking the ball by standing in its path, and Tkl – Tackles. Additionally, it exhibits relatively high values in Err – Errors and Def – Defensive actions that lead to a shot attempt compared to the other profiles. These characteristics are commonly associated with defenders.
  • Profile 3 displays remarkably high values in Tkl – Tackles, GCA – Goal-Creating Actions, Def – Defensive actions that lead to a shot attempt, PassDead – Completed dead-ball passes that lead to a shot attempt, PassLive – Completed live-ball passes that lead to a shot attempt, SCA – Shot-Creating Actions, CrsPA – Crosses into Penalty Area, PPA – Passes into Penalty Area, PiFT – Passes into Final Third, KP – Key Passes, PrgP – Progressive Passes, PrgC – Progressive carries. Additionally, it exhibits high values in Ast – Number of assists. These characteristics, combined with the consistently high values across all these variables, suggest an association with full-backs and midfielders. These positions involve moving the ball around the field and are frequently positioned centrally during the game.
  • Profile 4 demonstrates remarkably high values in ShToSh – Shots that lead to another shot attempt, TO – Successful take-ons that lead to a shot attempt, SoT – Shots on Target, Sh – Shots Total, and Gls – Number of goals. These characteristics are typically associated with pure strikers.
Figure 3 represents the memberships of the archetypal players (very high membership score) for each profile and a selection of other players that exhibit an interesting profile memberships. The archetypal players are:
  • Goalkeepers: Luigi Sepe of Salernitana team, who has a membership 0.962 in Profile 1.
  • Defenders: Federico Baschirotto of Lecce, who has a membership 0.973 in Profile 2.
  • Full-backs and midfielders: Cristiano Biraghi of Fiorentina, who has a membership 0.963 in Profile 3.
  • Pure strikers: Victor Osimhen of Napoli, who has a membership 0.947 in Profile 4.
Among the other players displayed, Chris Smalling, primarily a defensive player for Roma, but he also demonstrates notable scoring ability, having netted several goals in the season under review. This dual aspect of his play is captured by our model, which assigns him a modest membership to the pure strikers profile, acknowledging this secondary role. Edin Dzeko’s well-rounded abilities as a complete centre forward, blending technique with game vision, are reflected in his affiliation to the midfielders’ profile. Paulo Dybala, one of Serie A’s most technically skilled and elegant second forwards, often takes on ball-running duties in the forward area. This blend of a striker’s and a midfielder’s responsibilities is captured in our model. Likewise, Henrikh Mkhitaryan, an unconventional and agile trequartista, showcases a hybrid skill set that encompasses both attacking prowess and defensive contributions. Lastly, Wilfried Singo, known for his physical strength and speed as a right full-back or outside midfielder, is characterized by his dynamic play, adept in both offensive drives and defensive recoveries. As illustrated in Figure 3, the partial membership model adeptly encapsulates these diverse player characteristics, demonstrating its efficacy in capturing the complex nature of football players’ roles and abilities.
Figure 3: Partial membership model. A graphical representation of the archetypal football players and a selection of other interesting players. The pie chart for each player shows their partial membership to the profiles.
Based on the findings presented in Figure 3, it is evident that the model effectively captures the playing positions of the football players and successfully highlights nuances in their playing styles. The profile membership values and pie charts for each player in the database are displayed in the supplementary files of this paper.
These results could have practical implications for coaches, talent scouts, team managers, and analysts. These stakeholders can utilize the findings to make informed decisions related to team strategy, talent acquisition, and statistical research, ultimately enhancing performance and understanding in the field of football.
3.2 Mixed Membership Model Application
Mixed membership models are applied to count data in White and Murphy (2016). A detailed description of the model is given therein.
The profiles means for the mixed membership (MM) model applied to the football player performance data are presented in Table 5 and Figure 4. It must be emphasised, that in this application, the MM model fails in capturing the cited above archetypes. The players with highest membership per profile are: the striker Ciro Immobile with 0.591 in profile 1, the defensive player Gianluca Mancini with 0.485 membership in profile 2, the defensive player Martin Erlic with 0.700 membership in profile 3, the midfielder Jerdy Schouten with 0.533 membership in profile 4 and the second-tier Christian Kouamé with 0.548 membership in profile 5.
Table 5: Mixed membership model estimated expected profile means.
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 7.34 3.91 64.14 26.44 35.62 16.83 20.35 109.46
2 1.25 1.04 19.66 73.55 55.81 5.89 14.09 10.32
3 0.12 0.11 7.86 0.19 0.25 0.31 0.93 42.31
4 0.74 0.90 0.13 116.27 9.17 2.81 6.34 66.60
5 2.35 2.36 37.30 48.66 20.00 7.80 37.36 25.83
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 31.13 1.60 44.59 1.21 0.45 5.79 5.90 0.50
2 10.64 3.68 30.02 22.25 20.73 0.98 1.76 0.72
3 1.01 0.11 10.24 8.11 0.52 0.04 0.17 0.17
4 4.77 1.64 1.99 32.39 1.09 0.36 0.93 0.76
5 18.92 11.59 72.56 49.05 5.04 2.25 3.12 0.96
GCA Tkl Blocks Int Clr Err
1 8.77 10.51 10.01 4.51 6.17 0.22
2 2.42 0.96 17.60 29.90 16.28 0.33
3 0.35 22.30 0.10 17.71 84.78 1.17
4 2.35 33.45 31.19 15.49 20.77 0.64
5 5.28 50.20 19.57 0.36 46.82 0.27
Figure 4: Mixed membership model. Expected profiles means, conditional on profile membership, with 5 profiles.
The interpretation of the profile means in the MM model is far to be immediate compared to the one of PM model.
  • Profile 1 suggests a grouping of strikers, but it also exhibits high mean values in variables more typical of midfielders, such as PrgC – Progressive carries.
  • Profile 2 shows high mean values in Int – Interceptions, Blocks – Number of times blocking the ball by standing in its path, PassDead – Completed dead-ball passes that lead to a shot attempt, and Sh – Shots Total. These characteristics could be associated with defensive midfielders, although Sh - Shots Total is not typical for this role.
  • Profile 3 generally has low values in most variables, indicating a potential grouping of goalkeepers. However, it also presents high values in Int — Interceptions and Tkl – Tackles, which are more commonly associated with defenders.
  • Profile 4 exhibits very high values in PrgP – Progressive Passes, Blocks – Number of times blocking the ball by standing in its path, and Tkl – Tackles. This profile could be interpreted as grouping offensive midfielders, even though Blocks are not typically associated with this role.
  • Profile 5 might group defenders and full-backs, but these two positions have distinct characteristics, making the interpretation somewhat challenging.
The membership values resulted by the MM model and the respective pie charts for each player in the database are displayed in the supplementary files of this paper.
3.3 Finite Mixture Model Application
In a similar context of this paper, mixture models are applied to count data in White and Murphy (2016) to compare the results with those achieved with the mixed membership model. For a more in-depth explanation of the mixture modelling framework please refer to Everitt and Hand (1981) and Bouveyron et al (2019).
We fitted the multivariate Poisson mixture model to the player data and the profiles means are presented in Table 5 and Figure 4. In this application, similar to the MM model, the mixture model encounters challenges in effectively capturing the archetypal players. It’s important to clarify, as highlighted in the Introduction (see Section 1), that the finite mixture models function differently compared to mixed and partial membership models. Specifically, in finite mixture models, the membership of each unit to the clusters is binary, rather than continuous between 0 and 1, as is the case in mixed and partial membership models. This distinction necessitates a different interpretation of the results. In finite mixture models, the posterior probability of cluster membership should be viewed as reflecting uncertainty rather than partial membership. Nonetheless, for a comprehensive understanding, we include the results derived from the finite mixture model.
The players with highest membership per profile are: the attacking midfielder Filip Duricić with 0.570 in profile 1, the full-back Giovanni Di Lorenzo with 0.574 membership in profile 2, the midfielder Filip Kostić with 0.620 membership in profile 3, the defensive player Danilo Luiz da Silva with 0.605 membership in profile 4, the midfielder Stefano Sensi with 0.609 membership in profile 5 and the full-back Rogério Oliveira da Silva with 0.609 membership in profile 6.
Table 6: MM model expected profiles means
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 7.85 2.94 29.21 120.45 41.50 18.24 40.96 41.33
2 2.98 3.44 46.24 78.69 24.83 9.03 25.05 23.75
3 0.75 1.29 7.79 55.00 9.23 3.49 15.50 103.26
4 0.08 0.14 72.16 0.19 64.54 0.09 0.35 11.53
5 0.51 0.27 0.14 37.22 0.22 1.72 3.22 62.41
6 0.77 16.43 18.52 11.94 3.58 8.15 1.37 5.72
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 13.31 2.10 1.75 32.32 0.31 4.69 6.00 0.55
2 28.90 12.49 70.33 50.62 5.83 3.65 3.47 0.96
3 14.42 4.53 39.42 29.15 1.39 1.10 1.37 0.69
4 0.50 0.09 3.83 2.47 0.69 0.09 0.09 0.14
5 2.55 0.37 10.63 8.56 21.38 0.11 0.47 0.41
6 1.52 21.68 16.23 0.68 0.23 1.38 0.91 2.06
GCA Tkl Blocks Int Clr Err
1 6.41 10.27 9.51 3.62 17.89 0.21
2 7.78 24.28 15.94 11.20 5.41 0.20
3 2.95 34.89 22.23 18.89 33.38 0.42
4 0.21 2.95 0.16 0.48 54.07 1.00
5 1.05 0.90 18.81 19.66 89.60 1.00
6 53.16 32.30 30.60 13.01 0.59 0.05
Figure 5: Finite mixture model. Expected profiles means, conditional on profile membership, with 6 profiles.
It is clear that in the mixture model, the interpretation of the profile means is less immediate than the PM model or even the MM model.
3.4 Model Comparison
When comparing clustering models, one of the key parameters to consider is the interpretability of the results. This aspect is extensively discussed in Fraley and Raftery (1998) and Forgy (1965). The ability to interpret the clusters and derive meaningful insights from them is crucial in various domains, including sports analysis.
When comparing the models, it becomes evident that the interpretation of the clusters is significantly easier and more accurate in the PM model compared to the MM model. This disparity can be attributed to the inherent differences in the underlying assumptions of the two models.
The PM model, unlike the MM model, does not assume factorization over attributes. This means that each data attribute of a given data point is not assumed to be drawn independently from a mixture distribution based on the membership vector. In contrast, MM models are designed to handle situations where the objects being modeled consist of exchangeable sub-objects.
The lack of such assumptions in the PM model enhances its interpretability in our application. It allows for a more straightforward and intuitive understanding of the clusters, as the model does not impose strict dependencies between the attributes. As a result, the PM model excels in capturing the nuances of different playing positions without forcing the interpretation to conform to specific attribute relationships. The ease and quality of cluster interpretation are superior in the PM model compared to the MM model and even more when compared with the mixture model. The latter, with its crisp clustering approach, falls short of meeting the specific objectives and requirements of our proposed application. Additionally, the PM model demonstrates superior capability in identifying archetypal players, a key aspect in our analysis.
Overall, the PM model’s flexibility, underpinned by its independence from attribute-related assumptions, enables it to more accurately represent the diverse playing positions in football. This is achieved without the burden of imposing restrictive assumptions on the dataset, making it a highly suitable choice for this application.
4 Conclusions And Future Developments
Partial membership models offer analysts significantly greater flexibility compared to traditional model-based clustering or standard distance-based clustering methods. In our study, we tailored the model specifically for count data and applied it to the analysis of Serie A football players during the 2022/2023 season. Our goal was to estimate the various positions players tend to occupy on the field, in addition to their primary positions, by analyzing their playing styles. We based our analysis on a set of 22 count variables recorded during games, carefully selected to cover the key skills pertinent to each player’s role. This application also allowed us to test the model’s reliability by comparing its results with each player’s actual playing position. Furthermore, when compared with mixed membership and mixture models for count data, the partial membership model yielded more realistic and interpretable results. It excelled particularly in capturing archetypal players, whose distinct characteristics define specific clusters. These archetypes not only aid in clarifying each cluster’s definition but also contribute to the model’s overall stability and interpretability. By serving as benchmarks, they assist in better identifying and understanding the roles of players whose classifications might otherwise be ambiguous. While the partial membership model does require considerable computational resources, we believe its potential applications extend beyond sports analysis, offering valuable insights in fields like social sciences, genetics, natural sciences, and textual analysis. It can address some limitations inherent in mixed membership models. In the realm of sports analysis, our findings could significantly benefit coaches, talent scouts, team managers, and analysts. Utilizing these insights, they can make more informed decisions regarding team strategy, talent acquisition, and statistical research, thereby enhancing both performance and understanding in football.
For future developments, two primary areas present intriguing opportunities for enhancing our model. Firstly, addressing the issue of over-dispersion, which is a common challenge in count data, is of considerable interest. Developing methods to accurately assess and incorporate over-dispersion into the model would enable more precise and reliable predictions, particularly in datasets where variance significantly exceeds the mean. Secondly, adding a temporal dimension to the model opens up another avenue for exploration. By incorporating a temporal aspect, the model could provide a dynamic view of the clusters and allow for more nuanced analyses.
Acknowledgements
Thomas Brendan Murphy’s research was supported by the Science Foundation Ireland Insight Research Centre (12/RC/2289_P2) and a visiting period at Collegium de Lyon.
Appendix A Application: Washington DC Bike Data
We apply partial membership model on the data of the bike sharing company of Washington DC. The data are collected daily, from June 15th to July 15th, 2022, and record each single ride: date and time of start of trip, date and time of end of trip, name, ID, longitude and latitude of starting station, name, ID, longitude and latitude of ending station222 The data are freely available at https://capitalbikeshare.com/system-data. Figure 6 shows the number of trips per station in the considered time period.
Figure 6: Number of bicycle trips per bike sharing station from 15 of June to 15 of July 2022
We calculated the number of times bikes are collected from each of the 660 stations and we modelled these counts using a partial membership model, with the intent to explore the interactions between the bikes stations usage, to improve the allocation of the bikes. Partial membership model suits this type of application because the bikes move between the stations along the day, so the stations usage could vary and their membership could be partial. It should be noted that we do not addressed the temporal dependency as the temporal nature of the data would require. Nevertheless, the approaches appear to identify interesting behaviour in the data, and serve to illustrate the usefulness of the method. We run the model over a range of $K=1,\dots,6$. The model with the lowest WAIC is the one with 5 profiles (or components). For a better visualization, in Figure 7 are represented the natural log of the profiles means, while Figure 8 shows the marginal simplices representing stations’ profile membership.
Figure 7: Log of the expected number of rides per day from June 15th to July 15th, 2022, conditional on profile membership, with 5 profiles.
Figure 8: Marginal simplices representing bikes stations’ profile membership
Figure 9 represents as a pie chart the profile membership for each of the bike stations.
Figure 9: Bike stations’ pie charts of profiles membership.
It could be seen that Profile 5 groups the busiest stations, which are mainly located in the center of the city. Profile 1 the less used ones, which are mainly in the outlying areas, Profile 2 is an average usage stations cluster and looking at the map, it seems to connect the centre to the peripheral areas, Profile 3 groups the stations mostly used during the weekends, with an high peak of usage during the holiday of July 4th (Monday), which is public holiday in the USA. The stations with an high membership to this profile, are often located near the river or green ares, or also in the outlying areas. Profile 4 is the group of the stations mostly used on working days.